OpenStudy (anonymous):

Find the Integration of e^x *x^2

OpenStudy (anonymous):

You can use integration by parts: \[\int\limits f(x)g'(x)=f(x)g(x)-\int\limits f'(x)g(x)\]

OpenStudy (anonymous):

Lets identify f(x) and g(x) here: because e^x integral is easy it will be g(x) and x^2 is f(x) You will need to use this method once again to finish it all Answer will be e^x x^2 -e^x 2x +e^x 2=e^x(x^2-2x+2)